OpenStudy (anonymous):

Solve for B. A = 3(B + C) B = (A - C)/3 B = (3A - C)/3 B = (A - 3C)/3

OpenStudy (anonymous):

something's wrong with the q

OpenStudy (anonymous):

B=(A-3C)/3 Multiply the denominator up and it works.

OpenStudy (anonymous):

i think u've provided the options only and not the expression from which we r supposed to find the answer

OpenStudy (anonymous):

B=(A/3) - C B=(A/3) - (3C/3) B=(A-3C)/3

OpenStudy (anonymous):

B = (3(B+C) - C) / 3 B = (3B + 2C) / 3 (3(B+C) - C) = (3(3(B+C)) - C) = (3(B+C) - 3C) 2B = (3(3(B+C)) - C) + (3(B+C) - 3C) 2B = 12B + 8C 8C = 10B C = 10B/8 -10B = 8C B = 8C/10 B = (4 / 5) / -19 B = 4/-95

OpenStudy (anonymous):

The question might be missing a few parts, and in that case the answer would be wrong. But if that's the question, then my answer should be correct.

OpenStudy (anonymous):

I agree, I just noticed it isn't clear what the original equation was or if one has been presented with 4 equations to solve for B.